# Fast Polynomial Root-Finding and Approximation of the Eigenvalues of a Matrix

### v36

@@ 23.2 The case where a disc contains all $d$ roots
INPUT: a positive tolerance $\epsilon=1/2^{b}\leq 1$ and a black box polynomial $p(x)$ of a degree $d$ such that
@@
OUTPUT: a complex $z$, and a positive $r$ such that #$\#(D(z,r))=d$ and either $r\leq\epsilon$ or otherwise the disc $D(z,r)$ is $\frac{2}{13}$-rigid.
@@
(ii) Compute an approximation $\tilde{r}\geq 0$ to $r_{1}(z)$ by applying Alg. 9.1 for $m=d$, a constant $\beta>1$, and $\epsilon$ replaced by $\epsilon/4$.
If $\tilde{r}\leq 3\epsilon/4$, output $z$ and $r=\epsilon$, defining $\epsilon$-compression into the disc $D(z,r)$ of radius $\epsilon$.
Otherwise output $r=\tilde{r}+0.25\epsilon$, defining compression into rigid disc $D(z,r)$.
Correctness of the algorithm is readily verified if Alg. 9.1 outputs a range $[\rho_{-},\rho_{+}]$ of length at most $\epsilon/4$, for in this case $r_{1}\leq r=\tilde{r}+\epsilon/4\leq\epsilon$, and #$\#(D(z,\epsilon))=d$.
Otherwise $1/\beta\leq\tilde{r}/r_{1}\leq\beta$ for a constant $\beta\geq 1$ of Alg. 9.1, while the disc $D(x,r_{1})$ is $\frac{1}{2\alpha+2}$-rigid by virtue of Cor. 23.1 where $\alpha\leq 2.25$ by virtue of THm. 23.1 because $y\geq 4$ in our case. It follows that the disc $\bar{D}:=D(z,r_{1}(z))$ is $\frac{2}{13}$-rigid.
@@
By the same argument we can ensure that unless $\tilde{r}\leq 3\epsilon/4$ the disc $D(z,r_{1}(z))$ is $\eta$-isolated for any $\eta>4$ if we initialize Alg. 23.1 at a sufficiently large $y$.
@@
This compression queries evaluation oracle twice while performing Schröder’s single iteration for computing $z$ and otherwise runs at the cost of performing Alg. 9.1, estimated in Cor. 9.1. We can decrease the latter cost bound for a large class of polynomials $p$ by choosing larger tolerance $\epsilon$. Then we will yield desired compression if the algorithm outputs $r$ exceeding $\epsilon/4$. Otherwise we can recursively repeat the computations for tolerance recursively decreased to $\epsilon/4$.
23.3 Extension to any well-isolated disc
@@
Now suppose that #$\#(D(0,1))=m<d$ and $i(D(0,1))\geq\theta$ for a sufficiently large $\theta$ of order $d^{O(1)}$. Then verify that the $d-m$ external roots, lying outside the disc $D(0,1)$, little affect the proof of Thm. 23.1 and our whole study in the previous subsection.
Thus we first perform $O(\log(d))$ subdivision steps (see Observation 20.2). We do not need compression if these steps break the component into more than one component or compress it into an $\epsilon$-disc. Otherwise the component remains compact and hence in $O(\log(d))$ subdivision steps shrinks enough to become $d^{\nu}$-isolated for any fixed constant $\nu$, in which case we apply our compression algorithm.
23.4 A fast root-counter and a fast e/i test based on Schröder’s iterations
Let the unit disc be sufficiently strongly isolated and let a single Schröder’s iteration (23.1) be applied at a point $y$ lying sufficiently far from the unit disc $D(0,1)$.
Then our study in this subsection implies that the iteration outputs a point $z=z(m)$ lying in or near the disc $D(0,1)$, say, in the disc $D(0,2)$, if and only if #$\#(D(0,1))=m$, that is, we count roots in $D(0,1)$ at the cost of evaluation of $p(x)$ at two points.
Indeed, if #$\#(D(0,1))=m$, then, as we have showed in this section, the iteration outputs a point $z=z(m)$ lying near the convex hull of the root set in $D(0,1)$ and hence lying in or near the disc $D(0,1)$, say, in the disc $D(0,2)$ if the value $|y|$ is large enough.
In this case $|m~{}{\rm NR}(y)-y|\leq 2$, and so $|{\rm NR}(y)-y|/m\leq 2/m$.
If #$\#(D(0,1))=\bar{m}\neq m$, then the iteration outputs a point $z=z(m)$ that lies at the distance at least $|\bar{m}-m|\cdot|{\rm NR}(y)|-2$ from the disc $D(0,2)$, which must be large if $|y|$ is sufficiently (although reasonably) large.
Extension to a fast e/i test. Now suppose that the disc $D(0,1)$ is sufficiently well isolated and apply this root-counter to the polynomial $xp(x)$ for a large $y$ and $m=1$, in which case Schröder’s iterations turn into Newton’s. In this case root-counter will show that the disc $D(0,1)$ contains no roots of $p(x)$ if and only if its output lies in that disc.
23.5 Overall complexity of our root-finders
We have reduced our root-finding in an isolated disc to performing at most $12m-$ 1 compression steps and in addition $O(m)$ e/i tests. We have also reduced a compression step to (i) performing $O(\log(d))$ additional subdivision steps (included to strengthen isolation of input discs for compression), (ii) application of Schröder’s single iteration, which evaluates polynomial $p(x)$ at two points, and (iii) performing $O(\log(b))$ times $m$-tests for approximation of the root radius $r_{d-m+1}(z,p)$.
Overall our root-finders apply e/i tests $O(m\log(d))$ times, $m$-tests $O(m\log(b))$ times, and at most $22m-$ 2 Schröder’s iterations. Clearly, the computational cost of Schröder’s is dominated, while in Sec. 15 we reduced an $m$-test is to $O(1)$ e/i tests, specified in Sec. 14. Based on the latter results we can strengthen our estimates of Sec. 18 and arrive at the complexity bounds of our Main Theorems. Alternatively, instead of estimating the cost of e/i tests and reduction to them of $m$- tests, we can use our estimates for the complexity of $\ell$- tests of Part II for $\ell=1$ and $\ell=m$ and would still arrive at our Main Theorems 2 – 4.
@@ 23.5 Overall complexity of our root-finders
Part V: Factorization, Root Isolation, and Functional Iterations
24 Polynomial Factorization, its links to Root-finding, and some bit-complexity estimates
24.1 Polynomial Factorization and its links to Root-finding (the results)
@@
The root-finders of [134, 96, 103] solve Problems 0 and 0 by extending the the solution of the following problem of independent importance:
Problem F. Approximate Polynomial Factorization: [^35] The solution of Problem F can be also extended to isolation of the zeros of a square-free polynomial with integer coefficients and approximation of matrix eigenvalues (see our Remark 25.4 and Sec. 11) and has various other applications to modern computations, e.g., to time series analysis, Wiener filtering, noise variance estimation, co-variance matrix computation, and the study of multichannel systems [150, 18, 8, 35, 36, 147]. For a polynomial of degree $d$ the known sharp upper and lower bounds on the input precision and number of bit operations for root-finding exceed those for numerical factorization by a factor of $d$. Given a real $b$ and $d+1$ coefficients $p_{0},p_{1},\dots,p_{d}$ of (1.1), compute $2d$ complex numbers $u_{j},v_{j}$ for $j=1,\dots,d$ such that
@@
[96, 97, 99, 100, 103] proved the following near optimal bound for Problem F and readily extended it to near optimal solution of Problem 0.
@@
In Sec. 24.3 we reduce Problem 0 to Problem F and vice versa by applying Schönhage’s results of [134, Cor. 4.3 and Sec. 19]:
@@
(i) $\mathbb{B}_{0}(b,d)=\tilde{O}(\mathbb{B}_{F}(bd,d))$ for any polynomial $p(x)$ of degree $d$, (ii) $\mathbb{B}_{F}(b,d)=\tilde{O}(\mathbb{B}_{0}(b+d+\log(d),d))$ for any polynomial $p(x)$ of degree $d$, and (iii) $\mathbb{B}_{0}(b,d)=\tilde{O}(\mathbb{B}_{F}2(b,d))$ for polynomials of Def. 24.1, with bounded cluster sizes.
By combining Thms. 24.1 and 24.2 obtain the following estimates.
@@
One can solve Problem 0 by using $\tilde{O}((b+d)d+(b+m)m^{2})$ bit operation for $m$ of Def. 1.1 of Sec. 1.5. The bound turns into $\tilde{O}((b+d)d$ if $m^{2}=\tilde{O}(d)$.
@@
(ii) apply the algorithm that supports Thm. 24.3 to the $m$th degree factor of the polynomial $\prod_{j=1}^{d}(u_{j}x-v_{j})$ of (24.1) that shares with this polynomial all its zeros lying in the disc $D(1+\beta)$ for positive $\beta$ of Def. 1.1.363636We can slightly speed up stage (i) because at that stage we actually only need to approximate factorization of the factor $f$ of degree $m$ of $p$ that shares with $p$ its zeros in $D$.
24.2 Lower bounds on the bit operation complexity

### v37

@@ 17.4 Compression algorithm for a disc containing $d$ roots
INPUT: a positive $\epsilon=1/2^{b}\leq 1$ and a black box polynomial $p(x)$ of a degree $d$ such that
@@
OUTPUT: a complex $z$, and a positive $r$ such that #$\#(D(z,r))=d$ and the disc $D(z,r)$ is $\frac{2}{13}$-rigid unless $r\leq\epsilon$.
@@
(ii) Compute an approximation $\tilde{r}\geq 0$ to $r_{1}(z)$ by applying Alg. 16.1 for $\epsilon$ replaced by $\epsilon/4$.
Output $z$ and choose $r$ to define $\epsilon$-compression into the disc $D(z,r)$.
Namely, output $r=\epsilon$ if $\tilde{r}\leq 3\epsilon/4$.
Otherwise output $r=\tilde{r}+0.25\epsilon$, then the disc $D(z,r)$ is rigid.
Correctness of the algorithm is readily verified if Alg. 16.1 outputs a range $[\rho_{-},\rho_{+}]$ of length at most $\epsilon/4$, for in this case $r_{1}\leq r=\tilde{r}+\epsilon/4\leq\epsilon$, and #$\#(D(z,\epsilon))=d$.
Otherwise $1/\beta\leq\tilde{r}/r_{1}\leq\beta$ for a constant $\beta\geq 1$ of Alg. 16.1, while the disc $D(x,r_{1})$ is $\frac{1}{2\alpha+2}$-rigid by virtue of Cor. 17.2 where $\alpha\leq 2.25$ by virtue of Thm. 17.1 because $y\geq 4$ in our case. It follows that the disc $\bar{D}:=D(z,r_{1}(z))$ is $\frac{2}{13}$-rigid.
@@
If we initialize Alg. 17.1 at a sufficiently large $y$, then the same argument ensures that the disc $D(z,r_{1}(z))$ is $\eta$-isolated for any $\eta>4$ unless $\tilde{r}\leq 3\epsilon/4$.
@@
This compression queries evaluation oracle twice while performing Schröder’s single iteration for computing $z$ and otherwise runs at the cost of performing Alg. 16.1, estimated in Cor. 16.1.
We decrease the latter cost for a large class of polynomials $p$ by choosing a larger tolerance $\epsilon$. Then we yield desired compression if the algorithm outputs $r$ exceeding $\epsilon/4$. Otherwise we recursively repeat the computations for the tolerance recursively decreasing to $\epsilon/4$.
17.5 Compression of an isolated disc
@@
If $m<d$ and $i(D)\geq d^{\nu}$ for a sufficiently large constant $\nu$, then Eqn. (3.1) implies that the $d-m$ external roots, lying outside the disc $D$, little affect the proof of Thm. 17.1 and our whole study in the previous two subsections as long as we apply Schröder’s iteration at a complex point $y$ lying far enough from both disc $D$ and the circle $C(0,d^{\nu})$.
The MCD of a compact component can be isolated but not strongly. Then, however, we apply subdivision steps to that component and in order of $\log(d)$ steps have it broken or strongly isolated by virtue of Observation 17.2. Hence we can extend our study in the previous subsections. $O(\log(d))$ subdivision steps for compact components involve $O(m\log(d))$ e/i tests. The factor $\log(d)$ is immaterial for the $\tilde{O}$ estimates of our Main Theorems, but we remove it next.
17.6 From isolation to strong isolation
Let $s_{1}$ denote the sum of all $m$ roots in $D$. Clearly, the point $s_{1}/m$ lies in MCD and even in the convex hull of the root set $X(D)$. We compute approximation $c=s_{1}(b^{\prime})/m$ of $s_{1}/m$ within $1/2^{b^{\prime}}$ at the cost of $O(b^{\prime})$ queries of an evaluation oracle (see Thm. 12.1). For $b^{\prime}$ of order $\log(d)$ we yield error bound $1/d^{\nu}$ for a fixed positive constant $\nu$ of our choice at the cost of $O(\log(d))$ queries of an evaluation oracle 12.1).
By applying Alg. 16.1 we compute approximation $\rho$ of $r_{d-m+1}(c)$ also within $1/d^{\nu}$, and so
$|c-s_{1}/m|\leq 1/d^{\nu}~{}{\rm and}~{}|\rho-r_{d-m+1}(c)|\leq 1/d^{\nu}.$ (17.5)
The rigidity $\eta=\eta(D(c,\rho))$ of the disc $D(c,\rho)$ is given by the value $\frac{1}{\rho}\max^{*}|x_{i}-x_{j}|$ where maximum is over all pairs of roots lying in the disc $D(c,\rho)$. Now notice that
$\max^{*}|x_{i}-x_{j}|>0.5~{}r_{d-m+1}(s_{1}/m).$ (17.6)
Indeed, because of invariance of the claim under map (1.4), we can assume that $s_{1}=0$ and $x_{i}<0$. Then $\Re(x_{j})>0$ for some $j$, and (17.6) follows.
Furthermore, $|r_{d-m+1}(s_{1}/m)-r_{d-m+1}(c)|\leq|s_{1}/m-c|$; combine this bound with Eqn. (17.5) and obtain that $|r_{d-m+1}(s_{1}/m)-\rho|\leq 2/d^{\nu}$. Together with (17.6) this estimate implies that
$\max^{*}|x_{i}-x_{j}|>0.5(\rho-2/d^{\nu}).$
If $\rho\geq 4/d^{\nu}$, then $\rho-2/d^{\nu}\geq\rho/2$, the disc $D(c,\rho)$ is 0.25-rigid, and compression is completed.
If $\rho\leq 4/d^{\nu}$, then the disc $D(c,\rho)$, sharing with the unit disc $D(0,1))$ all its $m$ roots, is strongly isolated, namely, is $d^{\nu}/4$-isolated; then we can apply to it the compression algorithms of the previous sections.
18 Reduction of $m$-test to e/i tests
To complete our compression algorithms we should supply a $\sigma$-soft $\ell$-test for the input of Alg. 16.1, for which we can apply the low cost Las Vegas $\ell$-test in Sec. 13.
In application to compression, however, we can apply two alternative tests because we only need either (i) a $\sigma$-soft $m$-test for a sub-disc $D(0,\rho)$ of the unit disc $D(0,1)$ containing $m$ roots where $\rho<1$ or alternatively (ii) e/i test for the annulus $A(0,\rho,1)$ of Sec. 13.4.
(i) A $\sigma$-soft $m$-test for a sub-disc. Clearly, we can immediately reduce $d$-test in the unit disc $D(0,1)$ for $p(x)$ to performing 1-test (that is, e/i test) for that disc and for the reverse polynomial $p_{\rm rev}(x)$ of (1.5), but next we reduce $m$-test for any $m$, $1\leq m\leq d$, and for any disc $D$ to bounded number of e/i tests for sub-discs. If at least one of the e/i tests fails, then definitely there is a root in a narrow annulus about the boundary circle of an input disc, and then the internal disc of the annulus would not contain $m$ roots. Otherwise we obtain a root-free concentric annulus about the boundary circle of an input disc and then reduce $m$-test to root-counting in an isolated disc, which we can perform at a dominated cost (see claim (ii) of Prop. 12.1).
Algorithm 18.1.
Reduction of $m$-test to e/i tests.
INPUT: a polynomial $p(x)$ of a degree $d$, two values $\bar{\theta}>1$ and $\bar{\sigma}>1$, an integer $m$ such that $1\leq m\leq d$ and #$\#(D(0,\bar{\sigma}))\leq m$, and two algorithms $\mathbb{ALG}~{}0$ and $\mathbb{ALG}~{}1$, which for a black box polynomial perform $\bar{\sigma}$-soft e/i test for any disc and root-counting in an isolated disc, respectively.
$v=\lceil 2\pi/\phi\rceil$ (18.1)
$c_{j}=\exp(\phi_{j}\sqrt{-1}),~{}j=0,1,\dots,v,$ (18.2)
$c_{1}$$c_{2}$$c_{3}$$c_{4}$$c_{5}$$c_{6}$$c_{7}$$c_{8}$$c_{9}$$c_{0}$
Figure 7: $v=9$, $|c_{j}-c_{j-1}|=\rho>|c_{9}-c_{0}|$ for $j=1,2,\dots,9$.
To prove correctness of Alg. 18.1 we need the following result.
Theorem 18.1.
Proof.
$c_{0}$$c_{1}$$z$$z^{\prime}$
Let us prove that $|z_{j}-z_{j}^{\prime}|=\rho\sqrt{3}$ for $j=1$ and hence also for $j=2,\dots,v$.
Corollary 18.1.
Correctness of Alg. 18.1. (i) If algorithm $\mathbb{ALG}~{}1$ outputs inclusion in at least one of its applications, then one of the discs $D(c_{j},\bar{\sigma}\rho)$ contains a root, and then Alg. 18.1 correctly reports that $r_{d-m+1}\geq 1-\sigma\bar{\rho}$.
(ii) Otherwise Cor. 18.1 holds, and then correctness of Alg. 18.1 follows from correctness of $\mathbb{ALG}~{}1$.
Computational complexity. Alg. 18.1 invokes algorithm $\mathbb{ALG}~{}0$ at most $v+1$ times for $v=O(1)$ of (18.1). If $\mathbb{ALG}~{}0$ outputs exclusion $v+1$ times, then algorithm $\mathbb{ALG}~{}1$ is invoked and runs at a dominated cost.
Remark 18.1.
One can modify Alg. 18.1 as follows (see [48]): instead of fixing a positive $\rho$ and then computing $v$ of (18.1) fix an integer $v>0$, apply algorithm $\mathbb{ALG}~{}0$ at $v$ equally spaced points $c_{j}=\exp(\frac{2j\pi\sqrt{-1}}{v}),~{}j=0,1,\dots,v-1$, let $\phi=\frac{2\pi}{v}$, and then define $\rho$. We also arrive at this algorithm if we compute or guess this value $\rho$ and then apply Alg. 18.1 but notice that $c_{v}=c_{0}$ for $c_{j}$ of (18.2) and only invoke algorithm $\mathbb{ALG}~{}0$ for $c_{0},\dots,c_{v-1}$.
19 Counting e/i tests
19.1 An upper estimate
The upper estimate $22m-$ 2 of part (ii) of Sec. 4 on the number of compression steps is sharp – reached in the case of the complete binary tree $\mathbb{T}$.
Besides compression steps our subdivision root-finders involve e/i tests. Let us count them. Each compression step is preceded by an e/i test, and so we may need to perform at least $22m-$ 2 e/i tests. According to the main theorem of this section below, this lower bound is sharp up to a constant factor:
Theorem 19.1.
19.2 The first lemma
We begin our proof of Thm. 19.1 with
Lemma 19.1.
Proof.
[…]
19.3 The second lemma
Lemma 19.2.
Fix two integers $i\geq 0$ and $g\geq 3$, fix a value $\sigma$ of softness, $1<\sigma<\sqrt{2}$, and define the separation fraction $a:=2-\sigma\sqrt{2}$ of Observation 17.1. For all $j$ let $j$-th subdivision step process $v_{j}$ components made up of $w_{j}$ suspect squares, $n_{j}$ of which are non-adjacent, let $\Delta_{j}$ denote their side length, and let the $j$-th steps perform no compression for $j=i,i+1,\dots,i+g$. Then
Proof.
Now we only need to bound by $O(m)$ the number of e/i tests in the remaining subdivision steps, not interrupted with compression steps.
After a compression step we resume subdivision process with a suspect square whose side length bound decreases in the next two subdivision steps from at most $4\Delta_{j}$ to at most $\Delta_{j}$.
19.4 Counting e/i tests at partition steps
Let $h$ denote the height of our component tree and count suspect squares processed in subdivision steps that partition components, temporarily ignoring the other steps.
Notice that
$\sum_{i=0}^{h-g}v_{i+g}\leq\sum_{i=0}^{h}v_{i}-\sum_{i=0}^{g-1}v_{i}<2m-g,$
and so
[…]
Obtain that
Hence
19.5 Counting all e/i steps
Next consider a non-compact component, count all suspect squares processed in all subdivision steps applied until it is partitioned, and then prove in Lemma 19.3 that up to a constant factor the overall number of these suspect squares is bounded by their overall number at all partition steps. Such a result will extend to all subdivision steps the bound $n^{\prime}=O(m)$ of (19.3), so far restricted to partition steps. Together with Lemma 19.1 this will complete our proof of Thm. 19.1.
Lemma 19.3.
Let $k+1$ successive subdivision steps with $\sigma$-soft e/i tests be applied to a non-compact component for $\sigma$ of our choice in the range $1<\sigma<\sqrt{2}$. Let the $j$-th step output $n_{j}$ non-adjacent suspect squares for $j=0,1,\dots,k$. Then $\sum_{j=0}^{k-1}n_{j}\leq n_{k}\frac{10\sqrt{2}}{a}$ for the separation fraction $a:=2-\sigma\sqrt{2}$ (cf. Observation 17.1.)
Proof.
Hence $5n_{k}\geq n_{i}a2^{k-i-1.5}$ for all $i$, and so
Remark 19.1.
20 Overall complexity of our root-finders
We reduced our root-finding in a square with $m$ roots in its close neighborhood to $O(m)$ e/i tests in addition to at most $22m-$ 2 compression steps. We reduced compression first of an isolated and then of strongly isolated disc to (i) counting roots in an isolated disc ($O(\log(d))$ evaluation queries) (ii) application of Schröder’s single iteration (two evaluation queries), and (iii) approximation of the root radius $r_{d-m+1}(z,p)$, that is, performing $O(\log(b))$ $m$-tests.
In Sec. 18 we reduced an $m$-test to $O(1)$ e/i tests, and so the overall complexity of subdivision iterations is dominated by the complexity of performing $O(m\log(b))$ e/i tests and $O(m\log(d))$ queries of an evaluation oracle. By combining the latter estimates with those of Sec. 7 for an e/i test, we complete the proof of Main Theorems 1 – 4.
21 Algorithmic extras
Next we accelerate our root-finders for a large class of inputs.
1. To compress the isolated unit disc $D:=D(0,1)$ containing a small number $m$ of roots $z_{1},\dots,z_{m}$, compute $s_{1},\dots,s_{m}$ the first $m$ powers sums of the roots lying in $D$, then recover the coefficients of the factor $f(x)=\prod_{j=1}^{m}(x-z_{j})$ from Newton’s identities (cf. [83, Eqn. (2.5.4)]), approximate its zeros, and apply a black box root-finder to the polynomial $p(x)/f(x)$ represented by the pair of polynomials $p(x)$ and $f(x)$.
2. In the case of a large integer $m$ the latter algorithm suffers from swelling the coefficients of the factor $f(x)$. For $m=d$ the algorithm [102] empirically approximates all $m$ roots at a rather low cost, which provides more than just compression of $D$. Can we extend this approach to compression in the case where $m<d$ and the disc $D$ is sufficiently strongly isolated?
This is a research challenge, but we propose a similar low cost heuristic recipe which, as we conjecture, accelerates our root-finding for a large class of inputs.
Let the disc $D=D(0,1)$ be strongly isolated and let $i(D)\geq d^{2\nu}$ for a reasonably large $\nu$, and then instead of Newton’s apply Schröder’s single iteration (17.1) at sufficiently many, say 100, equally spaced points $y$ of the circle $C:=C(0,d^{\nu})$. Let $Z$ denote the set of images $z=y-m{\rm NR}(y)$. By extending our study in Sec. 17.5, deduce that all these images lie in the $\alpha\Delta$-neighborhood of the root set $X(D)$ for $\Delta:=\Delta(D)$ and $\alpha$ fast converging to 1 as $\nu$ grows large and hence that $\Delta(Z)\leq(2\alpha+1)\Delta$.
We may run into worst case inputs where $Z$ is a singleton and $\Delta(Z)=0$, but we conjecture that under a random rotation of the set of points $y$, the ratio $\Delta(Z)/\Delta$ exceeds a positive constant whp, and then we can narrow the initial range of BoL in Sec. 16 from $[-b,0]$ to $[\log_{2}(\Delta(Z)),0]$ at a low cost of performing Schröder’s iteration (17.1) at reasonably many, say 100, points.
@@ 21 Algorithmic extras
Part V: Factorization, Root Isolation, Functional Iterations, and Root Radii Estimation
We organize this part of the paper as follows. In Secs. 22 and 23 we cover some results by Schönhage [106] on polynomial factorization and its links to root-finding and root-isolation, and our lower and upper bit-complexity estimates for these problems. In Sec. 24 we recall root-finders by means of Newton’s, Weierstrass’s [119] (aka Durand-Kerner’s [23, 51]), and Ehrlich’s of 1967 [28], aka Aberth’s of 1973 [2] iterations, accelerate them by means of incorporation of FMM, and initialize Newton’s iterations for a black box input polynomial. In Sec. 25 we recall fast approximation of all $d$ root radii where the coefficients of $p$ are given.
22 Polynomial Factorization, Root-finding and Root Isolation:
Linking three Problems and their bit operation complexity
22.1 Linking Factorization and Root-finding: upper bounds on the bit operation complexity
@@
The root-finders of [106, 78, 84] solve Problems 0 and 0 by extending the solution of the following problem of independent importance:
Problem F. Approximate Polynomial Factorization: [^28] The solution of Problem F can be also extended to isolation of the zeros of a square-free polynomial with integer coefficients (see our Remark 23.4) and has various other applications to modern computations, listed in Sec. 1.1. For a polynomial of degree $d$ the known sharp upper and lower bounds on the input precision and number of bit operations for root-finding exceed those for numerical factorization by a factor of $d$. Given a real $b$ and $d+1$ coefficients $p_{0},p_{1},\dots,p_{d}$ of (1.1), compute $2d$ complex numbers $u_{j},v_{j}$ for $j=1,\dots,d$ such that
@@
In [78, 84] we proved the following near-optimal bound for Problem F.
@@
In Sec. 22.3 we reduce Problem 0 to Problem F and vice versa by applying Schönhage’s results of [106, Cor. 4.3 and Sec. 19]:
@@
(i) $\mathbb{B}_{0}(b,d)=\tilde{O}(\mathbb{B}_{F}(bd,d))$ for any polynomial $p(x)$ of degree $d$, (ii) $\mathbb{B}_{F}(b,d)=\tilde{O}(\mathbb{B}_{0}(b+d+\log(d),d))$ for any polynomial $p(x)$ of degree $d$, and (iii) $\mathbb{B}_{0}(b,d)=\tilde{O}(\mathbb{B}_{F}2(b,d))$ for polynomials of Def. 22.1, with bounded cluster sizes.
By combining Thms. 22.1 and 22.2 obtain the following estimates.
@@
One can solve Problem 0 by using $\tilde{O}((b+d)d+(b+m)m^{2})$ bit operation for $m$ of Def. 1.1 of Sec. 1.5. The bound turns into $\tilde{O}((b+d)d)$ if $m^{2}=\tilde{O}(d)$.
@@
(ii) apply the algorithm that supports Thm. 22.3 to the $m$th degree factor of the polynomial $\prod_{j=1}^{d}(u_{j}x-v_{j})$ of (22.1) that shares with this polynomial all its zeros lying in the disc $D(1+\beta)$ for positive $\beta$ of Def. 1.1.292929We can slightly speed up stage (i) because at that stage we actually only need to approximate factorization of the factor $f$ of degree $m$ of $p$ that shares with $p$ its zeros in $D$.
22.2 Linking Factorization and Root-finding: lower bounds on the bit operation complexity